OpenStudy (jjuden):

Arrange the following fractions in order from least to greatest. 2/3, 3/5, 5/7

OpenStudy (anonymous):

2/3 = 70/105 3/5 = 63/105 5/7 = 75/105

OpenStudy (anonymous):

3/5, 2/3, 5/7